OpenStudy (anonymous):

1/7(14-7 p) - 2=-2(1/2p+3)+6 solve the equation step by step please

OpenStudy (anonymous):

lets start with \[\frac{1}{7}\left(14-7p\right)\] what do you get when you multiply out using the distributive law?

OpenStudy (anonymous):

2-1p ?

OpenStudy (anonymous):

yes, good

OpenStudy (anonymous):

now we have on the left \[2-p-2\] or just \(-p\)

OpenStudy (anonymous):

on the right, start with \[-2\left(\frac{1}{2}p+3\right)\] and multiply what do you get for this one?

OpenStudy (anonymous):

-1p-6?

OpenStudy (anonymous):

yup. so far so good now what is \(-p-6+6\)?

OpenStudy (anonymous):

-p-12?

OpenStudy (anonymous):

not quite what is \(-6+6\) or if you prefer \(6-6\)?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

right, so what is \(-p-6+6\)?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

well actually you are left with \(-p\)

OpenStudy (anonymous):

so on the left we removed parentheses, combined like terms, and ended up with \(-p\) and on the right the same thing happened, we also ended up with \(-p\) now the equation looks like \[-p=-p\]

OpenStudy (anonymous):

oh ok so thats the answer

OpenStudy (anonymous):

the actual answer is "this is an identity" since \(-p=-p\) is always true in other words, both the left and right hand side of this equation are always the same, not matter what you put in for \(p\)

OpenStudy (anonymous):

*no matter

OpenStudy (anonymous):

Oh ok thank you very much (:

OpenStudy (anonymous):

yw